The annual yield on government-issued bonds in Mexican Peso is 15%. The analyst says that Peso will depreciate by 10% in the following year. If the Fisher Effect holds, what should be the yield on the one-year US T-Bills? Please write your answer with three decimal places? For instance, 1.2% is 0.012. Answer:
Answer #1

As per fisher effect

F/S = (1+ra)/(1+rb)
F= forward rate
S = spot rate.
ra = interest rate of price currency. 15%
rb= interest rate of base currency.

If Mexican peso depreciates by 10% then

F/S = 1.1/1 = 1.1

1.1 = (1+0.15)/(1+rb)
rb= 4.54%

yield on the one-year US T-Bills is 4.54%